OpenStudy (kayders1997):

S''(t)=5+4t-2t^2 initial velocity v(0)=3 m/s initial displacement s(0(=10m find position after t seconds

OpenStudy (kayders1997):

@zepdrix @satellite73

OpenStudy (kayders1997):

@mathmale

OpenStudy (fortytherapper):

S''(t) is a(t), but they want v(t) and s(t)

OpenStudy (kayders1997):

Sorry it's A''(t) for beginning

OpenStudy (kayders1997):

Okay so anti differentiate

OpenStudy (fortytherapper):

Right, it will then equal v(t), since the derivative of v(t) is a(t)

OpenStudy (kayders1997):

Okayso V'(t)=5t+2t^2-t^3?

OpenStudy (snowsurf):

Since S''(t) = a(t) where a is the acceleration you need to integrate to get the velocity function.

OpenStudy (fortytherapper):

I would check it out Does taking the derivative of your v(t) give you their a(t)?

OpenStudy (snowsurf):

Yes \[\frac{ dv(t) }{ dt } = a(t)\]

OpenStudy (snowsurf):

So the velocity will be \[v(t) = \int\limits_{}^{} 5+4t -2t ^{2}\]

OpenStudy (fortytherapper):

\[v(t) = 5t-2t^2-t^3\] So a(t) = \[5-4t-3t^2\], which is different from theirs

OpenStudy (kayders1997):

Oh so I was wrong? :O oops

OpenStudy (kayders1997):

Would it be -2/3t^3 and plus I the middle term?

OpenStudy (fortytherapper):

Right, it would just be 5-4t-2/3t^3

OpenStudy (fortytherapper):

I mean 5t-2t^2-2/3t^3

OpenStudy (kayders1997):

Hmmmmm it's plus 4t in the middle?

OpenStudy (fortytherapper):

Mb, Im mixing numbers up lol

OpenStudy (kayders1997):

It's okay :)

OpenStudy (kayders1997):

Do it again?

OpenStudy (fortytherapper):

Right, let's do it now. But we forgot 1 thing The +C

OpenStudy (kayders1997):

Omg I always forget the c!!!!

OpenStudy (fortytherapper):

So it's \[v(t) = -\frac{ 2 }{ 3 } t^3-2t^2+5t+C\] They give that v(0) = 3. So let's plug in 0 for t

OpenStudy (kayders1997):

Okay

OpenStudy (fortytherapper):

\[\frac{ 2 }{ 3 }(0)^3 - 2(0)^+5(0)+C = 3\]

OpenStudy (fortytherapper):

Or something like that lol

OpenStudy (kayders1997):

So 0+c=3

OpenStudy (kayders1997):

C=3

OpenStudy (fortytherapper):

Right, so now we have \[v(t)=-\frac{ 2 }{ 3 }t^3+2t^2+5t+3\] They want s(t) also, so let's integrate again

OpenStudy (fortytherapper):

I have to go, but do the exact same thing we did for v(t), and that function is your answer Remember the +C

OpenStudy (kayders1997):

s(t)=ugh this is getting to hard blah!

OpenStudy (kayders1997):

Thank you so much! :)

OpenStudy (fortytherapper):

We can integrate this one together if you like

OpenStudy (kayders1997):

If you have to go it's okay

OpenStudy (fortytherapper):

Nah it's cool. It won't take long. Since it's adding/subtracting we can separate each part Whats the integral of -2/3t^3? 2t^2? 5t? 3? Which one of those can you do right now?

OpenStudy (kayders1997):

3 I can do right now, 3t

OpenStudy (fortytherapper):

Alright, that one is good For the rest, it's basically the power rule backwards So for 5t... thats also \[\frac{ 5t^1 }{ 1 } \] \[\frac{ 5t ^{1+1} }{1+1 }\]

OpenStudy (fortytherapper):

1 + 1 is the exponent on the t +1